x/3 = 4/x+1

the directions say to solve the equation

Answer by stanbon(48512)

x/3 = 4/(x+1)
x(x+1) = 12
x^2 + x - 12 = 0
(x+4)(x-3) = 0
x = -4 or x = 3
=====================
Cheers,
Stan H.